OpenStudy (anonymous):

A painting is purchased for $500. If the value of the painting doubles every 5 years, then its value is given by the function V(t) = 500 • 2t/5, where t is the number of years since it was purchased and V(t) is its value (in dollars) at that time. What is the value of the painting ten years after its purchase? $1,000 $1,400 $1,800 $2,000

OpenStudy (anonymous):

i really need help guys

OpenStudy (radar):

Maybe it would help if you saw the function as it should be written:\[V(t)=500\times 2^{(t/5)}\]Now plug in the value for t as stated in the problem. Then operate on the 2 exponently then multiply your result by 500.

OpenStudy (radar):

\[V(t) = 500\times 2^{(10/5)}\]

OpenStudy (anonymous):

so v(t)=2000

OpenStudy (radar):

\[V(t)=500\times 2^{2}\] Solve.

OpenStudy (radar):

Good work @mahmoud10

OpenStudy (radar):

Don't for the $ sign lol

OpenStudy (anonymous):

alright thanks man i really appreciate it

OpenStudy (radar):

You're welcome, good luck with your studies.

OpenStudy (anonymous):

thanks man